OpenStudy (anonymous):

complete the table for the equation y=-3x+5 and graph the equation x: -1 0 3 4 y:

OpenStudy (hoblos):

just replace the value of x in the equation to get the y value

razor99 (razor99):

hey mayliz whats up

OpenStudy (hoblos):

for -1 : y=-3(-1)+5 = 3+5 = 8 so the first point is (-1,8)

OpenStudy (hoblos):

do that for all the points then plot then and join

OpenStudy (mayliz):